OpenStudy (anonymous):

trigonometry finding exact identity value for cos 7pi/12?

OpenStudy (raden):

it will be easier if we convert to degrees form, first because pi=180 degrees cos ( 7pi/12) = cos (7*180/12) = cos(105) deg then, we know that 105 = 60+45, so cos105 = cos(45+60) now, use the formula : cos (a+b) = cosa*cosb -sina*sinb

OpenStudy (anonymous):

thankss!

OpenStudy (raden):

yw